OpenStudy (anonymous):

Evaluate the piecewise function for x = -1 and x = 4. h(x)= 2x + 1 if x < 2 x^2-4 if x >2

OpenStudy (anonymous):

sup sexy

OpenStudy (jennychan12):

h(-1) = 2(-1) +1 = 0 h(4) = (4)^2 -4 = 16-4=2